OpenStudy (inexilewetrust):

Piano movers hoist a 200 kg piano 10 m high in 30 s using a rope and simple pulley. What is their power output? 70 W 700 W 7,000 W 6 × 10^5 W

OpenStudy (shamim):

U know w=mgh=200*9.8*10=?

OpenStudy (shamim):

P=w/t=?

OpenStudy (inexilewetrust):

No, could you explain it a but more... easier?

OpenStudy (inexilewetrust):

*bit

OpenStudy (shamim):

Do u know p=w/t?

OpenStudy (shamim):

Here p=power w=work t=time

OpenStudy (shamim):

Work per unit time is called power

OpenStudy (shamim):

Response plz for better understanding!!!

OpenStudy (shamim):

U know w=F*s=mgh

OpenStudy (shamim):

Here F=force acting on object=mg s=distance travelledcby object=h

OpenStudy (shamim):

Given m=200 kg h=10m g=9.8m/s

OpenStudy (inexilewetrust):

w=F*s=mgh m=200 kg h=10m g=9.8m/s w=F*s=(200*10*9.8) w=F*s= 19600

OpenStudy (inexilewetrust):

Sorry, I was reading the formula and trying to see how on google.

OpenStudy (shamim):

Its ok

OpenStudy (inexilewetrust):

So I got 19600 as the answer.

OpenStudy (shamim):

Feel free to ask for more clarification!!"

OpenStudy (inexilewetrust):

Okay is 19600 the correct answer?

OpenStudy (shamim):

No

OpenStudy (shamim):

p=200*9.8*10/30=?

OpenStudy (inexilewetrust):

Okay is 19600 the correct answer?

OpenStudy (shamim):

p=653.3w

OpenStudy (inexilewetrust):

Can you show me how you got the answer?